OpenStudy (anonymous):

Find S25 for 1/2 + 1 + 3/2 + 2 +... A. 160 B. 162 C. 160.5 D. 162.5

OpenStudy (anonymous):

Do you know if its arithmetic or Geometric? Which do you think it is?

OpenStudy (anonymous):

If its Arithmetic, find d(common difference) If its Geometric, find r(common ratio)

OpenStudy (anonymous):

hmmm @chicagoblkhaks we cant proceed if you dont reply also lol

OpenStudy (anonymous):

sorry, im in a haystack of homework. but im not sure how to find these ones.

OpenStudy (anonymous):

1/2 + x = 1 Solve for x.

OpenStudy (anonymous):

x = 1/2

OpenStudy (anonymous):

1 + x = 3/2 Solve for x.

OpenStudy (anonymous):

3/2 + x = 2 Solve for x.

OpenStudy (anonymous):

Yes, x = 1/2

OpenStudy (anonymous):

So: First term in the sequence: 1/2 1/2 + 1/2 = 1 <--- Second term in the sequence 1 + 1/2 = 3/2 <--- Third term in the sequence 3/2 + 1/2 = 2 <---- Fourth term in the sequence

OpenStudy (anonymous):

okay so how do i keep that going to figure out the answer. cause my choices are all super high compare to that lol

OpenStudy (anonymous):

so is it a arithmetic sequence or geometric sequence? As you see, it increases by + 1/2

OpenStudy (anonymous):

i guess im not sure?

OpenStudy (anonymous):

is that considered a difference or a ratio/